OpenStudy (cheynicole2001):

Match the systems of equations with their solutions. solution: system of linear equations: 1. (-31, -56) A. x-y=25, 2x+5y=180 2. (51, 26) B. 2x-3y=-5, x+y/11=50 3. (47, 33) c. x-y=19, -2x+y/6=28 4. (-17, -36) d. 2x-y=-6, x-305/3=2y

OpenStudy (mtblue7):

1=d

OpenStudy (cheynicole2001):

thanks! do you happen to know of the others?

OpenStudy (ramirocruzo):

a. { x = 305/7, y = 130/7 } b. { x = 47, y = 33 } c. { x = -17, y = -36 } d. { x = -341/9, y = -628/9 } These are the solutions :D IDK the matches, just brute solved the equations

OpenStudy (cheynicole2001):

omg.. thank you soo mucchhh :)))

OpenStudy (ramirocruzo):

>~< No problemo...